# Spectrum of Schrödinger operators on subcovering graphs

### v2

@@ 2. Main results
2.1. Primitive sets in a lattice
Let $\Gamma$ be a lattice in ${\mathbb{R}}^{d}$ with a basis $\{\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}\}$. A set $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}\subset\Gamma$, where $d_{o}<d$, is said to be primitive if $\mathfrak{t}$ is a basis for the lattice $\Gamma\cap\mathop{\mathrm{span}}\nolimits(\mathfrak{t})$ or, equivalently, if $\mathfrak{t}$ can be completed to a basis of $\Gamma$. Here $\mathop{\mathrm{span}}\nolimits(\mathfrak{t})$ is the set of all linear combinations of the vectors from $\mathfrak{t}$ with real coefficients. When $d_{o}=1$, the single vector $\mathfrak{t}_{1}$ in $\mathfrak{t}$ is called a primitive vector. In this case we will often omit the subscript 1 in $\mathfrak{t}_{1}$ and just denote this single vector by $\mathfrak{t}$.
There is a useful characterization of primitive sets. Let ${\mathcal{T}}$ be the $d_{o}\times d$ matrix whose rows are the coordinates of $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$ with respect to the basis $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$. It is known, see, e.g., [C97], that $\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$ is a primitive set for the lattice $\Gamma$ if and only if the greatest common divisor of all the $d_{o}$-th order minors of the matrix ${\mathcal{T}}$ is one. In particular, a vector $\mathfrak{t}=t_{1}\mathfrak{a}_{1}+\ldots+t_{d}\mathfrak{a}_{d}$ of the lattice $\Gamma$ is primitive if and only if its components $t_{1},\ldots,t_{d}$ (with respect to the basis $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$ of $\Gamma$) are coprime integers, i.e., $\mathfrak{t}$ is not an integral multiple of other lattice vectors (other than ±$\pm\mathfrak{t}$). This characterization of primitive sets is independent of the choice of the basis of the lattice $\Gamma$, i.e., if the rows of the matrix ${\mathcal{T}}$ form a primitive set and $U$ is a $d\times d$ unimodular matrix (which relates different bases of the same lattice), then the rows of the matrix ${\mathcal{T}}U$ also form a primitive set.
@@ 2.1. Primitive sets in a lattice
i) Primitive vectors of a lattice are also called visible, since only they can be seen from the origin of the lattice. Non-primitive lattice points are hidden from view by other lattice points lying in the line of sight.
ii) Each vector in a primitive set of a lattice $\Gamma$ is a primitive vector. But not every set of primitive vectors in $\Gamma$ is a primitive set. For example, $\mathfrak{t}_{1}=(1,1,0)$ and $\mathfrak{t}_{2}=(1,-1,0)$ are primitive vectors of the lattice ${\mathbb{Z}}^{3}$, but $\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ is not a primitive set for ${\mathbb{Z}}^{3}$. Indeed, $\mathop{\mathrm{span}}\nolimits(\mathfrak{t}_{1},\mathfrak{t}_{2})={\mathbb{R}}^{2}\times\{0\}$, so
@@ 2.2. Spectra of Schrödinger operators on subcoverings with primitive sets of chiral vectors
Let ${\mathcal{G}}$ be a $\Gamma$-periodic graph with the fundamental graph ${\mathcal{G}}_{*}={\mathcal{G}}/\Gamma=({\mathcal{V}}_{*},{\mathcal{E}}_{*})$, where $\Gamma$ is a lattice of rank $d$. From now on we will consider only subcoverings ${\mathcal{G}}_{\mathfrak{t}}$ of ${\mathcal{G}}$ with primitive sets $\mathfrak{t}\subset\Gamma$ of chiral vectors $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$, $d_{o}<d$.
We describe the spectrum of the Schrödinger operators $H_{\mathfrak{t}}$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$. Denote by $K_{j}^{\pm}$ the level sets corresponding to the band edges $\lambda_{j}^{\pm}$ of the Schrödinger operator $H$ on ${\mathcal{G}}$, i.e.,
@@
Let ${\mathcal{G}}$ be a $\Gamma$-periodic graph, and $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$ be a primitive set of the lattice $\Gamma$. Then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ has the form
@@
for some $k_{o}\in K_{j}^{\pm}$, where $K_{j}^{\pm}$ are given by (2.1), and ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8), (1.9). In particular,
@@
ii) Let $\tau$ be the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$, where ${\mathcal{T}}^{\top}$ is the transpose of ${\mathcal{T}}$. Then
@@
ii) From Proposition 2.2 it also follows that a $\Gamma$-periodic graph ${\mathcal{G}}$ is isospectral to its subcovering ${\mathcal{G}}_{\mathfrak{t}}$, if and only if for each $j\in{\mathbb{N}}_{\nu}$ the condition (2.3) holds for some $k^{-}\in K_{j}^{-}$ and for some $k^{+}\in K_{j}^{+}$, where $K_{j}^{\pm}$ are given by (2.1). By isospectrality of periodic graphs we mean that the spectra of the Schrödinger operators on the graphs consist of the same number of bands and the corresponding bands coincide as sets.
iii) If the set $\mathfrak{t}\subset\Gamma$ is not primitive, then Proposition 2.2 is not true any more. For example, if $d_{o}=1$ and the single chiral vector $\mathfrak{t}=p\mathfrak{t}_{o}$ for some primitive $\mathfrak{t}_{o}\in\Gamma$ and integer $p\geqslant 2$, then the fundamental graph of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ has $p\nu$ vertices, and, consequently, the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on ${\mathcal{G}}_{\mathfrak{t}}$ consists of $p\nu$ spectral bands. Note that ${\mathcal{G}}_{\mathfrak{t}_{o}}$ is a $p$-fold subcovering of ${\mathcal{G}}_{p\mathfrak{t}_{o}}$, which yields that $\sigma(H_{\mathfrak{t}_{o}})\subseteq\sigma(H_{p\mathfrak{t}_{o}})$, see, e.g., [A95].
iv) Since $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$ are linearly independent, the matrix ${\mathcal{T}}{\mathcal{T}}^{\top}$ is positive definite. The condition $\tau\to\infty$ implies, in particular, that the Euclidean norm $|t_{s}|$ of each chiral index $t_{s}\in{\mathbb{Z}}^{d}$, see (1.8), tends to infinity. Indeed, the diagonal entries of the matrix ${\mathcal{T}}{\mathcal{T}}^{\top}$ are $|t_{1}|^{2},\ldots,|t_{d_{o}}|^{2}$. Then the smallest eigenvalue $\tau^{2}$ of ${\mathcal{T}}{\mathcal{T}}^{\top}$ satisfies
$\tau^{2}\leqslant\min\big{\{}|t_{1}|^{2},\ldots,|t_{d_{o}}|^{2}\big{\}},$
see, e.g., Theorem 4.3.15 in [HJ85]. This yields that $|t_{s}|\to\infty$, $s\in{\mathbb{N}}_{d_{o}}$, as $\tau\to\infty$. Note that if $d_{o}=1$, then $\tau=|t_{1}|$.
@@
vi) Due to the band structure of the spectrum of the Schrödinger operators on periodic graphs, there are some notions of isospectrality for periodic graphs. The problem of Floquet isospectrality for the discrete Schrödinger operators on the lattice ${\mathbb{Z}}^{d}$ was studied in [GKT93, K89, L23]. Fermi isospectrality on ${\mathbb{Z}}^{d}$ was discussed in [L24].
@@
Let $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$ be a primitive set of the lattice $\Gamma$, and let for some $j\in{\mathbb{N}}_{\nu}$ the band function $\lambda_{j}(\cdot)$ of the Schrödinger operator $H$ on a $\Gamma$-periodic graph ${\mathcal{G}}$ satisfy the Assumption A. Then the lower band edge $\lambda_{j}^{-}(\mathfrak{t})$ (the upper band edge $\lambda_{j}^{+}(\mathfrak{t})$) of the Schrödinger operator $H_{\mathfrak{t}}$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ of ${\mathcal{G}}$ has the following asymptotics
@@
Here $\mathrm{Hess}\,\lambda_{j}(k_{o})$ is the Hessian of $\lambda_{j}$ at $k_{o}$; ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8), (1.9); $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$, and ||$\cdot$|| ⋅ | denotes the Euclidean norm of a vector.
In particular, if $d_{o}=1$, i.e., the set $\mathfrak{t}$ consists of a single primitive vector with a chiral index $t\in{\mathbb{Z}}^{d}$, then
$\begin{array}[]{l}\displaystyle\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm}\mp\frac{1}{2}\cdot\frac{|x_{o}|^{2}}{|{\bf H}^{-1/2}\,t|^{2}}+g(\mathfrak{t}),\qquad\textrm{where}\\[12.0pt] x_{o}=(\langle t,k_{o}\rangle\mod 2\pi)\in(-\pi,\pi],\\[8.0pt] g(\mathfrak{t})=\textstyle O\big{(}\frac{1}{\tau^{3}}\big{)},\qquad\textrm{as}\qquad\tau=|t|\to\infty.\end{array}$ (2.6)
@@
i) The second term in the asymptotics (2.5) is equal to zero if and only if ${\mathcal{T}}k_{o}\in 2\pi{\mathbb{Z}}^{d_{o}}$. This agrees with Proposition 2.2, see the condition (2.3).
@@
iii) Although Assumption A1 is often used to establish many important properties of periodic operators (for example, the notion of electron’s effective masses in solid state physics [AM76]), for some periodic discrete graphs this assumption fails (see examples in [FK18, BK20]).
iv) For a particular case when $d_{o}=1$ and the single chiral index has the specific form $t=(t_{1},\ldots,t_{d-1},-1)\in{\mathbb{Z}}^{d}$, i.e., $t_{d}=-1$, the statements of Proposition 2.2 and Theorem 2.4 were obtained in [S24], see also Remark 1.7.
@@
In this section we illustrate the obtained results by some simple examples of periodic graphs and their subcoverings. The proofs of the examples are based on Proposition 2.2 and Theorem 2.4 and given in Section 4.
First we consider the hexagonal lattice ${\bf G}$ which is a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{2}$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2}$, see Fig. 1a. The fundamental graph ${\bf G}_{*}$ of ${\bf G}$ consists of two vertices $v_{1}$ and $v_{2}$ and three multiple edges connecting these vertices, see Fig. 1d. Let $H=\Delta+Q$ be the Schrödinger operator on ${\bf G}$ with a $\Gamma$-periodic potential $Q$. Without loss of generality (add a constant to $Q$ if necessary) we may assume that the potential $Q$ is given by
@@
Let ${\bf G}_{\mathfrak{t}}$ be a nanotube with a primitive chiral vector $\mathfrak{t}=t_{1}\mathfrak{a}_{2}+t_{2}\mathfrak{a}_{2}\in\Gamma$. Then the spectrum $\sigma(H_{\mathfrak{t}})$ of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on ${\bf G}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.7) has the form
@@
$\varepsilon(\mathfrak{t})=\frac{\pi^{2}}{6q(t_{1}^{2}+t_{1}t_{2}+t_{2}^{2})}+\textstyle O\big{(}\frac{1}{|t|^{3}}\big{)},\quad\textrm{as}\quad|t|\to\infty,\qquad t=(t_{1},t_{2})\in{\mathbb{Z}}^{2}.$
@@
i) For any primitive $\mathfrak{t}\in\Gamma$, the nanotube ${\bf G}_{\mathfrak{t}}$ is asymptotically isospectral to the hexagonal lattice ${\bf G}$ as $|t|\to\infty$. If $t_{1}-t_{2}\in 3{\mathbb{Z}}$, then the Schrödinger operators $H$ on ${\bf G}$ and $H_{\mathfrak{t}}$ on ${\bf G}_{\mathfrak{t}}$ have the same spectrum, i.e., ${\bf G}$ and ${\bf G}_{\mathfrak{t}}$ are just isospectral. If $\mathfrak{t}$ is not primitive, i.e., $\mathfrak{t}=p\mathfrak{t}_{o}$ for some primitive $\mathfrak{t}_{o}\in\Gamma$ and integer $p\geqslant 2$, then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on ${\bf G}_{\mathfrak{t}}$ consists of $2p$ spectral bands, some of which may be degenerate. But, due to (1.12), the inclusion $\sigma(H_{p\mathfrak{t}_{o}})\subseteq\sigma(H)$ still holds.
For example, if ${\bf G}_{\mathfrak{t}}$ is the zig-zag nanotube with the non-primitive vector $\mathfrak{t}=2\mathfrak{a}_{1}$ (see Fig. 1c), then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on ${\bf G}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.7) has the form

### v3

@@ 2. Main results
2.1. Primitive sets in ${\mathbb{Z}}^{d}$
A set $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}\subset{\mathbb{Z}}^{d}$, where $d_{o}<d$, is said to be primitive if $\mathfrak{t}$ is a basis for the lattice ${\mathbb{Z}}^{d}\cap\mathop{\mathrm{span}}\nolimits(\mathfrak{t})$ or, equivalently, if $\mathfrak{t}$ can be completed to a basis of ${\mathbb{Z}}^{d}$. Here $\mathop{\mathrm{span}}\nolimits(\mathfrak{t})$ is the set of all linear combinations of the vectors from $\mathfrak{t}$ with real coefficients. When $d_{o}=1$, the single vector $\mathfrak{t}_{1}$ in $\mathfrak{t}$ is called a primitive vector. In this case we will often omit the subscript 1 in $\mathfrak{t}_{1}$ and just denote this single vector by $\mathfrak{t}$.
There is a useful characterization of primitive sets. Let ${\mathcal{T}}$ be the $d_{o}\times d$ integer matrix whose rows are the vectors $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\in{\mathbb{Z}}^{d}$. It is known, see, e.g., [C97], that $\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}\subset{\mathbb{Z}}^{d}$ is a primitive set if and only if the greatest common divisor of all the $d_{o}$-th order minors of the matrix ${\mathcal{T}}$ is one. In particular, a vector $\mathfrak{t}=(t_{1},\ldots,t_{d})\in{\mathbb{Z}}^{d}$ is primitive if and only if its components $t_{1},\ldots,t_{d}$ are coprime integers, i.e., $\mathfrak{t}$ is not an integral multiple of other vectors (other than ±$\pm\mathfrak{t}$) of ${\mathbb{Z}}^{d}$. This characterization of primitive sets is independent of the choice of the basis of ${\mathbb{Z}}^{d}$, i.e., if the rows of the matrix ${\mathcal{T}}$ form a primitive set and $U$ is a $d\times d$ unimodular matrix (which relates different bases of ${\mathbb{Z}}^{d}$), then the rows of the matrix ${\mathcal{T}}U$ also form a primitive set.
@@ 2.1. Primitive sets in ${\mathbb{Z}}^{d}$
i) Primitive vectors of the lattice ${\mathbb{Z}}^{d}$ are also called visible, since only they can be seen from the origin of the lattice. Non-primitive lattice points are hidden from view by other lattice points lying in the line of sight.
ii) Each vector in a primitive set is a primitive vector. But not every set of primitive vectors is a primitive set. For example, $\mathfrak{t}_{1}=(1,1,0)$ and $\mathfrak{t}_{2}=(1,-1,0)$ are primitive vectors of ${\mathbb{Z}}^{3}$, but $\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ is not a primitive set in ${\mathbb{Z}}^{3}$. Indeed, $\mathop{\mathrm{span}}\nolimits(\mathfrak{t}_{1},\mathfrak{t}_{2})={\mathbb{R}}^{2}\times\{0\}$, so
@@ 2.2. Spectra of Schrödinger operators on subcoverings with primitive sets of chiral vectors
Let ${\mathcal{G}}$ be a ${\mathbb{Z}}^{d}$-periodic graph with the fundamental graph ${\mathcal{G}}_{*}={\mathcal{G}}/{\mathbb{Z}}^{d}=({\mathcal{V}}_{*},{\mathcal{E}}_{*})$. From now on we will consider only subcoverings ${\mathcal{G}}_{\mathfrak{t}}$ of ${\mathcal{G}}$ with primitive sets $\mathfrak{t}\subset{\mathbb{Z}}^{d}$ of chiral vectors $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$, $d_{o}<d$.
We describe the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$. Denote by $K_{j}^{\pm}$ the level sets corresponding to the band edges $\lambda_{j}^{\pm}$ of the Schrödinger operator $H$ on ${\mathcal{G}}$, i.e.,
@@
Let ${\mathcal{G}}$ be a ${\mathbb{Z}}^{d}$-periodic graph, and $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}\subset{\mathbb{Z}}^{d}$ be a primitive set. Then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ has the form
@@
for some $k_{o}\in K_{j}^{\pm}$, where $K_{j}^{\pm}$ are given by (2.1), and ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8). In particular,
@@
ii) Let $({\mathcal{G}}_{\mathfrak{t}})$ be a sequence of the subcoverings of ${\mathcal{G}}$ such that the smallest singular value $\tau$ of their $d_{o}\times d$ chiral matrices ${\mathcal{T}}$ goes to infinity. Then
@@
ii) From Proposition 2.2 it also follows that a periodic graph ${\mathcal{G}}$ is isospectral to its subcovering ${\mathcal{G}}_{\mathfrak{t}}$ with a primitive set $\mathfrak{t}$, if and only if for each $j\in{\mathbb{N}}_{\nu}$ the condition (2.3) holds for some $k^{-}\in K_{j}^{-}$ and for some $k^{+}\in K_{j}^{+}$, where $K_{j}^{\pm}$ are given by (2.1). By isospectrality of periodic graphs we mean that the spectra of the Schrödinger operators on the graphs consist of the same number of bands and the corresponding bands coincide as sets.
iii) If the set $\mathfrak{t}\subset{\mathbb{Z}}^{d}$ is not primitive, then Proposition 2.2 is not true any more. For example, if $d_{o}=1$ and the single chiral vector $\mathfrak{t}=p\mathfrak{t}_{o}$ for some primitive $\mathfrak{t}_{o}\in{\mathbb{Z}}^{d}$ and integer $p\geqslant 2$, then the fundamental graph of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ has $p\nu$ vertices, and, consequently, the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on ${\mathcal{G}}_{\mathfrak{t}}$ consists of $p\nu$ spectral bands. Note that ${\mathcal{G}}_{\mathfrak{t}_{o}}$ is a $p$-fold subcovering of ${\mathcal{G}}_{p\mathfrak{t}_{o}}$, which yields that $\sigma(H_{\mathfrak{t}_{o}})\subseteq\sigma(H_{p\mathfrak{t}_{o}})$, see, e.g., [A95].
iv) Since $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$ are linearly independent, the matrix ${\mathcal{T}}{\mathcal{T}}^{\top}$, where ${\mathcal{T}}^{\top}$ is the transpose of ${\mathcal{T}}$, is positive definite. The smallest singular value $\tau$ of ${\mathcal{T}}$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$. The condition $\tau\to\infty$ implies that the Euclidean norm $|\mathfrak{t}_{s}|$ of each chiral vector $\mathfrak{t}_{s}\in{\mathbb{Z}}^{d}$ tends to infinity. Indeed, the diagonal entries of the matrix ${\mathcal{T}}{\mathcal{T}}^{\top}$ are $|\mathfrak{t}_{1}|^{2},\ldots,|\mathfrak{t}_{d_{o}}|^{2}$. Then the smallest eigenvalue $\tau^{2}$ of ${\mathcal{T}}{\mathcal{T}}^{\top}$ satisfies
$\tau^{2}\leqslant\min\big{\{}|\mathfrak{t}_{1}|^{2},\ldots,|\mathfrak{t}_{d_{o}}|^{2}\big{\}},$
see, e.g., Theorem 4.3.15 in [HJ85]. This yields that
$|\mathfrak{t}_{s}|\to\infty,\qquad s\in{\mathbb{N}}_{d_{o}},\qquad\textrm{as}\qquad\tau\to\infty.$ (2.5)
Note that if $d_{o}=1$, then $\tau=|\mathfrak{t}_{1}|$. The condition (2.5) in turn implies that when $\tau\to\infty$
$\bullet$ the ”radius” of the cylinder ${\mathbb{R}}^{d}/\Gamma_{\mathfrak{t}}$, in which the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ is naturally embedded (see Section 1.2), tends to infinity (recall that $\Gamma_{\mathfrak{t}}$ is a sublattice of ${\mathbb{Z}}^{d}$ with the basis $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$);
$\bullet$ the set ${\mathcal{B}}_{\mathfrak{t}}$ defined by (1.9) is dense in the Brillouin zone ${\mathcal{B}}=(-\pi,\pi]^{d}$.
@@
vi) Since the spectrum of the Schrödinger operators on periodic graphs has band structure, there exist some notions of isospectrality for periodic graphs. The problem of Floquet and Fermi isospectrality for the discrete Schrödinger operators on the lattice ${\mathbb{Z}}^{d}$ was studied in [GKT93, K89, L23] and [L24], respectively.
@@
Let $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}\subset{\mathbb{Z}}^{d}$ be a primitive set, and let for some $j\in{\mathbb{N}}_{\nu}$ the band function $\lambda_{j}(\cdot)$ of the Schrödinger operator $H$ on a ${\mathbb{Z}}^{d}$-periodic graph ${\mathcal{G}}$ satisfy the Assumption A. Then the lower band edge $\lambda_{j}^{-}(\mathfrak{t})$ (the upper band edge $\lambda_{j}^{+}(\mathfrak{t})$) of the Schrödinger operator $H_{\mathfrak{t}}$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ of ${\mathcal{G}}$ has the following asymptotics
@@
Here $\mathrm{Hess}\,\lambda_{j}(k_{o})$ is the Hessian of $\lambda_{j}$ at $k_{o}$; ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8); $\tau$ is the smallest singular value of ${\mathcal{T}}$, and ||$\cdot$|| ⋅ | denotes the Euclidean norm of a vector.
In particular, if ${\mathcal{G}}_{\mathfrak{t}}$ is a subcovering of ${\mathcal{G}}$ with a single primitive vector $\mathfrak{t}\in{\mathbb{Z}}^{d}$, then
$\begin{array}[]{l}\displaystyle\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm}\mp\frac{1}{2}\cdot\frac{|x_{o}|^{2}}{|{\bf H}^{-1/2}\,\mathfrak{t}|^{2}}+g(\mathfrak{t}),\qquad\textrm{where}\\[12.0pt] x_{o}=(\langle\mathfrak{t},k_{o}\rangle\mod 2\pi)\in(-\pi,\pi],\\[8.0pt] g(\mathfrak{t})=\textstyle O\big{(}\frac{1}{\tau^{3}}\big{)},\qquad\textrm{as}\qquad\tau=|\mathfrak{t}|\to\infty.\end{array}$ (2.7)
@@
i) Assumption A1 is often used to establish many important properties of periodic operators such as electron’s effective masses in solid state physics [AM76], Green’s function asymptotics [KKR17] and other issues. It is an old and widely believed conjecture in mathematical physics that generically (with respect to perturbations of the periodic potential and edge weights) Assumption A1 holds true. For recent progress in proving the band edges nondegeneracy conjecture see [FS24] and references therein.
@@
iii) The second term in the asymptotics (2.6) is equal to zero if and only if ${\mathcal{T}}k_{o}\in 2\pi{\mathbb{Z}}^{d_{o}}$. This agrees with Proposition 2.2, see the condition (2.3).
iv) For $\tau$ large enough, the error term $g(\mathfrak{t})$ in (2.6) and (2.7) satisfies $|g(\mathfrak{t})|\leqslant C\tau^{-3}$, where $C$ is a positive constant depending on the Hessian of the band function $\lambda_{j}$ at the extrema point $k_{o}$ and the number $d_{o}$ of the chiral vectors but not depending on the chiral matrix ${\mathcal{T}}$ (i.e., not depending on the choice of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$).
v) For a particular case when $d_{o}=1$ and the single chiral vector has the specific form $\mathfrak{t}=(t_{1},\ldots,t_{d-1},-1)\in{\mathbb{Z}}^{d}$, i.e., $t_{d}=-1$, the statements of Proposition 2.2 and Theorem 2.4 were obtained in [S24], see also Remark 1.6.
@@
In this section we illustrate the obtained results by some simple examples of periodic graphs and their subcoverings. In all examples we will consider the Schrödinger operator $H$ defined by (1.1), (1.2) with the edge weight $\omega\equiv 1$. The proofs of the examples are based on Proposition 2.2 and Theorem 2.4 and given in Section 4.
First we consider the hexagonal lattice ${\bf G}$ which is a ${\mathbb{Z}}^{2}$-periodic graph with the periods $\mathfrak{a}_{1},\mathfrak{a}_{2}$, see Fig. 1a. The fundamental graph ${\bf G}_{*}$ of ${\bf G}$ consists of two vertices $v_{1}$ and $v_{2}$ and three multiple edges connecting these vertices, see Fig. 1d. Let $H=\Delta+Q$ be the Schrödinger operator on ${\bf G}$ with a ${\mathbb{Z}}^{2}$-periodic potential $Q$. Without loss of generality (add a constant to $Q$ if necessary) we may assume that the potential $Q$ is given by
@@
Let ${\bf G}_{\mathfrak{t}}$ be a nanotube with a primitive chiral vector $\mathfrak{t}=(t_{1},t_{2})\in{\mathbb{Z}}^{2}$. Then the spectrum $\sigma(H_{\mathfrak{t}})$ of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on ${\bf G}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.8) has the form
@@
$\varepsilon(\mathfrak{t})=\frac{\pi^{2}}{6q(t_{1}^{2}+t_{1}t_{2}+t_{2}^{2})}+\textstyle O\big{(}\frac{1}{|\mathfrak{t}|^{3}}\big{)},\qquad\textrm{as}\qquad|\mathfrak{t}|\to\infty.$
@@
i) For any primitive $\mathfrak{t}\in{\mathbb{Z}}^{2}$, the nanotube ${\bf G}_{\mathfrak{t}}$ is asymptotically isospectral to the hexagonal lattice ${\bf G}$ as $|\mathfrak{t}|\to\infty$. If $t_{1}-t_{2}\in 3{\mathbb{Z}}$, then the Schrödinger operators $H$ on ${\bf G}$ and $H_{\mathfrak{t}}$ on ${\bf G}_{\mathfrak{t}}$ have the same spectrum, i.e., ${\bf G}$ and ${\bf G}_{\mathfrak{t}}$ are just isospectral. If $\mathfrak{t}$ is not primitive, i.e., $\mathfrak{t}=p\mathfrak{t}_{o}$ for some primitive $\mathfrak{t}_{o}\in{\mathbb{Z}}^{2}$ and integer $p\geqslant 2$, then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on ${\bf G}_{\mathfrak{t}}$ consists of $2p$ spectral bands, some of which may be degenerate. But, due to (1.11), the inclusion $\sigma(H_{p\mathfrak{t}_{o}})\subseteq\sigma(H)$ still holds.
For example, if ${\bf G}_{\mathfrak{t}}$ is the zig-zag nanotube with the non-primitive chiral vector $\mathfrak{t}=(2,0)$ (see Fig. 1c), then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on ${\bf G}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.8) has the form